OpenStudy (anonymous):

Solve: 2(b – 8) > 12.

OpenStudy (anonymous):

2b - 16 > 12 ?

OpenStudy (anonymous):

A. b < 20 B. b > 6 C. b > 14 D. b > 20

OpenStudy (theviper):

I think C.

OpenStudy (anonymous):

first, treat your "greater than" sign as an equal sign: 2(b-8)=12 now solve for b: 2(b-8)-12=0 2b-16-12=0 2b=28 b=14 now just change your "=" back to ">" B>14 so as long as "B" is greater than 14, this equation will hold true :P

OpenStudy (theviper):

\[\Huge{\color{red}{Woah!\color{green}{☻}}}\] @harharf

OpenStudy (anonymous):

your welcome (: